# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

### v1

@@ Abstract
Audio-aware large language models (ALLMs) have recently made great strides in understanding and processing audio inputs. These models are typically adapted from text-based large language models (LLMs) through additional training on audio-related tasks. However, this adaptation process presents two major limitations. First, ALLMs often suffer from catastrophic forgetting, where important textual capabilities such as instruction-following are lost after training on audio data. In some cases, models may even hallucinate sounds that are not present in the input audio, raising concerns about their reliability. Second, achieving cross-modal alignment between audio and language typically relies on large collections of task-specific question–answer pairs for instruction tuning, making the process resource-intensive. To address these issues, we leverage the backbone LLMs from ALLMs to synthesize general-purpose caption-style alignment data. We refer to this process as bootstrapping audio-language alignment via synthetic data generation from backbone LLMs (BALSa). Building on BALSa, we introduce LISTEN (Learning to Identify Sounds Through Extended Negative Samples), a contrastive-like training method designed to improve ALLMs’ ability to differentiate between present and absent sounds. We further extend BALSa to multi-audio scenarios, where the model either explains the differences between audio inputs or produces a unified caption that describes them all, thereby enhancing audio-language alignment. Experimental results indicate that our method effectively mitigates audio hallucinations while reliably maintaining strong performance in audio understanding, reasoning benchmarks, and instruction-following skills. Moreover, incorporating multi-audio training further enhances the model’s comprehension and reasoning capabilities. Overall, BALSa offers an efficient and scalable approach to the development of ALLMs.
@@ I Introduction
To address both the challenge of catastrophic forgetting and the burden of data curation, we propose a distinct approach that leverages a backbone LLM to generate synthetic data for audio-language alignment. Here, the backbone LLM refers to the large language model that serves as the core of ALLM architectures. We term this approach BALSa, short for Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs. While prior works [14, 15, 16, 17, 18] have explored similar conceptual directions, we are the first to apply this approach to general audio tasks, offering advantages in simplicity, flexibility, and scalability. Instead of relying on manually crafted or task-specific question-answer pairs, often constructed using proprietary LLMs and complex generation rules, BALSa leverages the backbone LLM’s understanding of audio metadata like sound event tags to synthesize training data in a prompt-driven, automatic, and extensible manner. Users only need to design simple generation prompts like Repeat the audio or List non-existent sounds in the audio, allowing BALSa to flexibly generate diverse audio-language data tailored to different learning objectives with minimal human effort.
BALSa offers a flexible and generalizable pipeline that enables audio-language alignment by learning from the behavioral patterns of the backbone LLM. For instance, it can generate not only descriptive data, such as natural language descriptions of events occurring in the audio, but also contrastive data that distinguish between sounds that are present and those that are absent, simply by varying the generation prompt. This highlights BALSa’s potential as a general-purpose framework for synthesizing diverse alignment data across a wide range of audio understanding tasks. As part of this framework, we refer to the incorporation of synthetic negative samples as LISTEN  [^1] LISTEN was introduced in our previous work [19]. This paper extends that work with a unified training framework, multi-audio modeling, and additional evaluations and experiments. (Learning to Identify Sounds Through Extended Negative Samples) [19], a contrastive-like method that helps models distinguish between present and absent sounds. Building on the flexibility of BALSa to generate diverse types of alignment data, our findings indicate that synthetic data generated by backbone LLMs plays a crucial role in model training. In particular, incorporating negative synthetic samples through LISTEN effectively mitigates hallucinations, thereby enhancing model reliability.
@@
    We propose a data construction pipeline, BALSa (Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs), which synthesizes audio-language alignment data using a backbone LLM and simple prompts, eliminating the need for task-specific question-answer data collection. Our study explores and validates the effectiveness of this approach, in which the backbone LLM analyzes audio metadata such as sound event tags and generates descriptions guided by a flexibly designed prompt.
@@
    Under the BALSa framework, we introduce LISTEN, a contrastive-style training strategy that incorporates synthetic negative samples to help models distinguish between present and absent sounds, thereby reducing audio hallucination. Experimental results demonstrate that synthetic data generated by the backbone LLM plays a crucial role in model training, and the LISTEN method effectively mitigates audio hallucinations, enhancing model reliability in real-world applications.
@@ II Related Works
Large language models (LLMs) [20, 21, 22, 23, 24] can perform zero-shot learning through instruction tuning. This capability is expected to emerge when strong cross-modal alignment is achieved, particularly in scenarios where a text-based LLM is integrated with multi-modal encoders.
@@
Following this trend, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8] extend instruction tuning to audio-related tasks by combining audio encoders with language models and training them in a multi-stage process. Qwen-Audio [1] adopts a two-stage training strategy. In the first stage, multi-task pre-training is conducted using a manually curated dataset that includes tasks such as acoustic scene classification, sound event classification, and audio question answering. To unify the training format across tasks, Qwen-Audio introduces task-specific tags that indicate the nature of each task. This stage aims to build a foundational understanding of audio signals. In the second stage, Qwen-Audio performs supervised instruction tuning to align the model more closely with human intent. To create training data, proprietary LLMs [11, 25, 26] are used to generate question–answer pairs based on text metadata, while additional audio-dialogue datasets are constructed using human annotations. Through this process, Qwen-Audio enhances the model’s ability in both audio understanding and reasoning. Qwen2-Audio [2] follows a similar approach, focusing on collecting high-quality supervised fine-tuning data through meticulous curation. SALMONN [3] also employs the same two-stage pipeline. During instruction tuning, it prioritizes tasks based on their relevance and manually collects task-specific datasets, including those for audio question answering. Like Qwen-Audio, it leverages proprietary LLMs to generate question–answer pairs grounded in the captions of existing audio corpora. LTU [4] constructs a large-scale audio question answering dataset, covering both closed-ended and open-ended formats. The closed-ended set includes tasks such as sound classification, audio captioning, and temporal analysis. Questions are generated using proprietary LLMs, while answers follow a fixed structure derived from rule-based algorithms. For open-ended data, which is more expensive and labor-intensive to create manually, LTU uses a rule-based pipeline known as audio instruction generation. This method inputs metadata such as audio captions and sound event tags into carefully designed prompts, allowing proprietary LLMs to produce diverse and contextually appropriate audio QA pairs.
@@
On the other hand, a distinct strategy for developing audio-aware large language models focuses on leveraging the backbone LLM itself to generate audio–language alignment data. These alignment data include general response-style text conditioned on audio, captions that describe the audio given textual conditions, or continuations of audio content based on textual prompts. Several prior works [14, 27, 16, 18, 17] have adopted this approach. For instance, AudioChatLlama [14] and DiVA [18] use the backbone language model to generate responses conditioned on textual transcriptions. BLSP [16] further extends this idea by generating continuations based on transcriptions, while BLSP-Emo [17] and DeSTA2 [27] incorporate additional paralinguistic information like emotion to produce more expressive continuations and captions.
@@
Inspired by these methods and motivated by the lack of research applying this strategy to general audio settings, our work investigates how to bootstrap audio–language alignment through synthetic data generation using a backbone LLM. We further explore various extensions of this idea to evaluate its generality and effectiveness across different scenarios. For example, we exploit the flexibility and scalability of this framework to introduce negative samples and multi-audio scenario into the training process.
This distinction highlights the key difference between our work and prior studies [1, 2, 3, 4, 5]: instead of relying on externally generated task-specific question-answer pairs, we leverage a backbone LLM to generate general-purpose alignment data directly, enabling a more streamlined and self-contained approach to integrating audio with language models. This provides a practical way to train audio-aware LLMs without the need to manually collect or rely on proprietary models to generate a large amount of task-specific question-answer data.
@@ III-A Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs
The rationale behind generating audio-language pairs using the backbone LLM is to enable the model to acquire audio understanding capabilities solely through an audio modality adapter. This approach ensures that the audio-aware large language model aligns its comprehension of audio with the text-based metadata information understood by the backbone LLM, without requiring adjustments to the backbone LLM itself. In this work, we refer to this approach as Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs (BALSa).
@@
In Equation 1, [Begin of audio] and [End of audio] are expressed in natural language format, with the purpose of indicating to the model that the information between them represents the content from the audio. When training audio-aware large language models, the position of ${P}_{{seed}}$ will be replaced with corresponding audio representation. These generated responses to the input context are then used as our training targets. Actual examples are demonstrated in Table I.
@@ III-C Model Design and Training Approach
Figure 1 illustrates the architecture of our model based on BALSa, applied to single-audio and multi-audio scenarios, respectively. Our model consists of three main components: audio encoder, backbone large language model, and audio modality adapter.
@@ VI Result
Second, we explore an alternative approach by providing the model with a set of pre-collected questions source from OpenAQA [4] and generating corresponding answers using the BALSa pipeline. We refer to this variant as BALSa-OpenAQA. Experimental results indicate that BALSa-OpenAQA, in most cases, perform slightly worse than those trained with BALSa-generated positive and negative samples. However, the performance gap is not substantial overall. On the other hand, if we directly train the model using the original question-answer pairs provided in OpenAQA [4], rather than generating answers through the BALSa approach, the experimental results show that models trained with BALSa-OpenAQA consistently outperform those trained directly on the original OpenAQA dataset across all evaluation benchmarks. This finding further reinforces the importance of using the original backbone LLM for generating training data, highlighting its crucial role in enhancing model performance. Third, we explore an approach that utilizes pre-defined rule-based templates to construct paired training data. However, models trained with this method not only perform poorly across evaluation benchmarks but also exhibit a parroting effect, where responses frequently fail to follow instructions and instead generate only audio captions.
@@
Based on the experimental results, we report the following findings and observations: The BALSa framework enables the generation of audio-language alignment training data using only simple generation prompts, such as basic captioning-based instructions like Repeat the audio. If additional pre-collected question-answer data like OpenAQA [4] is available, it can also be processed through the BALSa pipeline to generate training data, with the pre-collected questions serving to increase the diversity of generation prompts. However, this approach requires additional data collection and preparation efforts. The results highlight that the key factor in the BALSa framework is the use of the backbone LLM for data generation. Furthermore, BALSa can be extended to multi-audio scenarios, leveraging explanations between two audio inputs to enhance the model’s audio understanding capabilities.
@@
Building on this, we extend BALSa to the multi-audio setting. Experimental results indicate that this extension leads to performance improvements across multiple evaluation benchmarks. By comparing the positive-negative setting in the single-audio scenario with the discrimination setting in the multi-audio scenario, we observe an 8% improvement in audio hallucination performance, representing an approximately 10% increase. On the fundamental audio question-answering benchmarks, our method achieves performance improvements of 8% and 15% on the EDANSA and NonSpeech tasks, respectively. On SAKURA, we observe a nearly 10% increase in overall accuracy, along with an 8% improvement specifically on multi-hop questions. In contrast, improvements in Audio Entailment are relatively smaller. Overall, we find that BALSa-MA, which enables the model to learn from two audio inputs simultaneously, enhances its basic audio understanding abilities, thereby benefiting more advanced reasoning tasks.
@@
VII Discussion
Prior work [1, 2, 3, 4, 5, 6, 7] has primarily focused on achieving audio-language alignment by collecting and constructing large-scale, task-specific question-answer datasets for model training. In contrast, we adopt a different strategy in BALSa by leveraging backbone LLMs to generate general captioning-based alignment data. Experimental results demonstrate that BALSa achieves comparable performance to previous methods and even outperforms them on certain benchmarks, while also maintaining a balance with fundamental instruction-following abilities. It effectively combines strong audio understanding capabilities with robust instruction-following performance, and shows reliable behavior in mitigating audio hallucinations, highlighting its trustworthiness in practical scenarios.
In audio-language alignment, the most fundamental objective is to train models to understand audio content. We achieve this through a captioning-based alignment strategy. Meanwhile, reasoning ability is inherently provided by the backbone LLM, as we do not modify any of its parameters. The motivation behind this approach is that, given a sufficiently capable text-based LLM as the foundation, it is only necessary to train a modality adaptation module to enable the model to understand general audio inputs without modifying the parameters of the original language model. Such a design also reduces computational demands, allowing the LLM to process audio information while preserving its extensive pre-trained textual and instruction-tuned capabilities. On the other hand, our experimental results further suggest that training audio understanding capabilities can also be achieved using question-answering pair data. However, the key factor remains that the backbone LLM must be used to generate corresponding answers, ensuring effective audio-language alignment.
VIII Conclusion
In this work, we introduce BALSa, a framework that uses backbone LLMs to generate synthetic data for alignment, reducing the reliance on large collections of task-specific QA datasets. By leveraging a captioning-based alignment strategy, BALSa achieves strong audio understanding and reasoning while preserving the instruction-following capabilities of large language models. Our experimental results demonstrate that BALSa achieves comparable or superior performance to previous approaches while improving data and computational efficiency. These findings also highlights the importance of utilizing the original backbone LLM for generating audio-language alignment data. In addition, by incorporating LISTEN (Learning to Identify Sounds Through Extended Negative Samples), BALSa effectively mitigates audio hallucinations, further reinforcing its robustness in real-world applications.
Beyond single-audio alignment, we extend BALSa to more complex audio understanding settings. In multi-audio scenarios, this extension further enhances modality alignment and leads to better performance on audio understanding and reasoning tasks. We further examine the role of discrimination-based training and observe that learning to handle multiple audio inputs alone is sufficient to improve overall understanding, without requiring explicit comparative objectives. Overall, BALSa provides a scalable and efficient pathway for developing and advancing audio-aware large language models.

### v2

@@ Abstract
Audio-aware large language models (ALLMs) have recently made great strides in understanding and processing audio inputs. These models are typically adapted from text-based large language models (LLMs) through additional training on audio-related tasks. However, this adaptation process presents two major limitations. First, ALLMs often suffer from catastrophic forgetting, where crucial textual capabilities like instruction-following are lost after training on audio data. In some cases, models may even hallucinate sounds that are not present in the input audio, raising concerns about reliability. Second, achieving cross-modal alignment between audio and language typically relies on large collections of task-specific question–answer pairs for instruction tuning, making it resource-intensive. To address these issues, previous works have leveraged the backbone LLMs to synthesize general-purpose, caption-style alignment data. In this paper, we propose a data generation framework that produces contrastive-like training data, designed to enhance ALLMs’ ability to differentiate between present and absent sounds. We further extend our approach to multi-audio scenarios, enabling the model to either explain differences between audio inputs or produce unified captions that describe all inputs, thereby enhancing audio-language alignment. We refer to the entire ALLM training framework as bootstrapping audio-language alignment via synthetic data generation from backbone LLMs (BALSa). Experimental results indicate that our method effectively mitigates audio hallucinations while reliably maintaining strong performance on audio understanding and reasoning benchmarks, as well as instruction-following skills. Moreover, incorporating multi-audio training further enhances the model’s comprehension and reasoning capabilities. Overall, BALSa offers an efficient and scalable approach to developing ALLMs.
@@ I Introduction
To address both the challenge of catastrophic forgetting and the burden of data curation, there is a growing trend toward leveraging a backbone LLM to generate synthetic data for audio-language alignment [14, 15, 16, 17, 18, 19]. Here, the backbone LLM refers to the large language model that serves as the initialisation of the core component of ALLM architectures. Instead of relying on manually crafted or task-specific question-answer pairs, which are often generated using proprietary LLMs and complex rules, these approaches leverage the backbone LLM to synthesize training data in a prompt-driven, automatic, and extensible manner. The developers only need to design simple generation prompts, and the backbone LLM can flexibly generate diverse data with minimal human effort. While the conceptual direction of using backbone LLMs for data generation has been widely explored in the speech domain, this article is the first to focus on extending the idea to general audio tasks, demonstrating that it also offers advantages in simplicity, flexibility, and scalability in the audio domain.
This paper represents the framework, BALSa, short for Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs. Different from previous works using backbone LLMs for synthetic data generation, BALSa generates not only descriptive data, such as natural language descriptions of events occurring in the audio, but also contrastive data, which distinguish sounds that are present from those that are absent by simply varying the generation prompt [^1] The incorporation of the synthetic negative samples was first proposed in our previous conference paper, LISTEN (Learning to Identify Sounds Through Extended Negative Samples) [20]. This article extends that conference paper with a unified training framework, multi-audio modeling, and additional evaluations and experiments.. Like previous work of synthetic data generation, BALSa leverages the backbone LLM’s understanding of audio metadata, such as sound event tags, to synthesize training data. Consequently, we only need the prompts like Repeat the audio or List non-existent sounds in the audio to efficiently generate diverse audio-language data tailored to various learning objectives. BALSa offers a flexible and generalizable pipeline that facilitates audio-language alignment by learning from the behavioral patterns of the backbone LLM. This highlights BALSa’s potential as a general-purpose framework for synthesizing diverse alignment data for a wide range of audio understanding tasks. Building upon the flexibility of BALSa in generating diverse alignment data, our findings indicate that incorporating negative synthetic samples effectively mitigates hallucinations, thereby enhancing model reliability.
@@
    We present BALSa (Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs). A critical feature of BALSa is a contrastive-style training strategy that incorporates synthetic negative samples, enabling models to distinguish clearly between present and absent sounds. Experimental results demonstrate that BALSa effectively mitigates audio hallucinations, thereby enhancing the model’s reliability in real-world applications.
@@
    Prior works [14, 15, 16, 17, 18, 19] have explored and validated the effectiveness of synthetic data generation using backbone LLMs. In these studies, the backbone LLM analyzes audio metadata and generates descriptions guided by designed prompts, but they focus primarily on the speech domain. This article is the first to apply this paradigm specifically to the non-speech audio domain. BALSa synthesizes audio-language alignment data using simple prompts from audio metadata, such as sound event tags, without requiring task-specific question-answer collections.
@@ II Related Works
There is already extensive research on audio and speech-related large language models. Due to space constraints, this paper discusses only the studies directly relevant to our work. For a comprehensive overview of the broader framework of audio and speech LLMs, readers are encouraged to consult the survey papers [21, 22, 23, 24, 25, 26, 27, 28].
@@
Large language models (LLMs) [29, 30, 31, 32, 33] can perform zero-shot learning through instruction tuning. This capability is expected to emerge when strong cross-modal alignment is achieved, particularly in scenarios where a text-based LLM is integrated with multi-modal encoders. Following this trend, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8, 34] extend instruction tuning to audio-related tasks by combining audio encoders with language models and training them in a multi-stage process.
@@
Qwen-Audio [1] adopts a two-stage training strategy. In the first stage, multi-task pre-training is conducted using a manually curated dataset that includes tasks such as acoustic scene classification, sound event classification, and audio question answering. To unify the training format across tasks, Qwen-Audio introduces task-specific tags that indicate the nature of each task. This stage aims to build a foundational understanding of audio signals. In the second stage, Qwen-Audio performs supervised instruction tuning to align the model more closely with human intent. To create training data, proprietary LLMs [11, 35, 36] are used to generate question–answer pairs based on text metadata, while additional audio-dialogue datasets are constructed using human annotations. Through this process, Qwen-Audio enhances the model’s ability in both audio understanding and reasoning. Qwen2-Audio [2] follows a similar approach, focusing on collecting high-quality supervised fine-tuning data through meticulous curation. SALMONN [3] also employs the same two-stage pipeline. During instruction tuning, it prioritizes tasks based on their relevance and manually collects task-specific datasets, including those for audio question answering. Like Qwen-Audio, it leverages proprietary LLMs to generate question–answer pairs grounded in the captions of existing audio corpora. LTU [4] constructs a large-scale audio question answering dataset, covering both closed-ended and open-ended formats. The closed-ended set includes tasks such as sound classification, audio captioning, and temporal analysis. Questions are generated using proprietary LLMs, while answers follow a fixed structure derived from rule-based algorithms. For open-ended data, which is more expensive and labor-intensive to create manually, LTU uses a rule-based pipeline known as audio instruction generation. This method inputs metadata such as audio captions and sound event tags into carefully designed prompts, allowing proprietary LLMs to produce diverse and contextually appropriate audio QA pairs.
@@
On the other hand, a distinct strategy for developing speech- or audio-aware large language models focuses on leveraging the backbone LLM itself to generate synthetic data for alignment. These alignment data include general response-style texts conditioned on speech/audio, captions describing speech/audio under textual conditions, or continuation responses of speech/audio content based on textual prompts. Several previous works [16, 14, 17, 19, 18] have adopted this approach. Unlike studies that rely on externally generated task-specific question-answer pairs [1, 2, 3, 4, 5], these works utilized a backbone language model to generate general-purpose alignment data directly, enabling a more streamlined and self-contained approach to integrating speech/audio with language models. DeSTA [15] found that the approach leveraged speech captioning to bridge the gap between speech and text modalities, while DeSTA2 [14] further discovered that generating captions using the backbone LLM was more efficient. AudioChatLlama [16] and DiVA [19] used the backbone language model to generate responses conditioned on textual transcriptions. BLSP [17] further extended this idea by generating continuations based on transcriptions, while BLSP-Emo [18] incorporated additional paralinguistic information like emotion to produce more expressive outputs. DeSTA2 [14] further extended the use of paralinguistic cues, such as emotion, accent, and gender, to enhance the expressiveness of generated speech captions.
@@
Inspired by these methods and motivated by the lack of research applying this strategy to general audio settings, our work investigates how to bootstrap audio–language alignment through synthetic data generation using a backbone LLM. We further explore various extensions of this idea to evaluate its generality and effectiveness across different scenarios. For example, we exploit the flexibility and scalability of this framework to introduce negative samples and multi-audio scenario into the training process. This article provides a practical way to effectively train audio-aware LLMs without the need to manually collect or rely on proprietary models to generate a large amount of task-specific question-answer data.
@@ III-A Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs
The rationale behind generating audio-language pairs using the backbone LLM is to enable the model to acquire audio understanding capabilities solely through an audio modality adapter. This approach ensures that the audio-aware large language model aligns its comprehension of audio with the text-based metadata information understood by the backbone LLM, without requiring adjustments to the backbone LLM itself.
@@
In Equation 1, [Begin of audio] and [End of audio] are expressed in natural language format, with the purpose of indicating to the model that the information between them represents the content from the audio. When training audio-aware large language models, the position of ${P}_{{seed}}$ will be replaced with corresponding audio representation. These generated responses to the input context are then used as our training targets. Actual examples are demonstrated in Table I. In this work, we refer to the whole data generation process in this subsection as Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs (BALSa).
@@ III-C Model Design and Training Approach
Figure 1 illustrates the overall framework. The model comprises three primary components: an audio encoder, a backbone large language model, and an audio modality adapter.
@@ VI-B Comparison of Different Data Generation Approaches
Second, we explore an alternative approach by providing the model with a set of pre-collected questions source from OpenAQA [4] and generating corresponding answers using the BALSa pipeline. We refer to this variant as BALSa-OpenAQA. Experimental results indicate that BALSa-OpenAQA, in most cases, perform slightly worse than those trained with BALSa-generated positive and negative samples. However, the performance gap is not substantial overall. On the other hand, if we directly train the model using the original question-answer pairs provided in OpenAQA [4], rather than generating answers through the BALSa approach, the experimental results show that models trained with BALSa-OpenAQA consistently outperform those trained directly on the original OpenAQA dataset across all evaluation benchmarks. Third, we explore an approach that utilizes pre-defined rule-based templates to construct paired training data. However, models trained with this method not only perform poorly across evaluation benchmarks but also exhibit a parroting effect, where responses frequently fail to follow instructions and instead generate only audio captions.
@@
Based on the experimental results, we report the following findings and observations: The BALSa framework can generate audio-language alignment training data using only simple generation prompts, such as basic captioning-based instructions like Repeat the audio. If additional pre-collected question-answer data like OpenAQA [4] is available, it can also be processed through the BALSa pipeline to generate training data, with the pre-collected questions serving to increase the diversity of generation prompts. However, this approach requires additional data collection and preparation efforts. The finding in this article further reinforces the established practice within the community of using the original backbone LLM for generating training data [14], underscoring its essential role in improving model performance.
VI-C Performance of BALSa-MA
@@
Furthermore, BALSa can be extended to multi-audio scenarios, leveraging explanations between two audio inputs to enhance the model’s audio understanding capabilities. Building on this, we extend BALSa to the multi-audio setting (BALSa-MA). Experimental results indicate that this extension leads to performance improvements across multiple evaluation benchmarks. By comparing the positive-negative setting in the single-audio scenario with the discrimination setting in the multi-audio scenario, we observe an 8% improvement in audio hallucination performance, representing an approximately 10% increase. On the fundamental audio question-answering benchmarks, our method achieves performance improvements of 8% and 15% on the EDANSA and NonSpeech tasks, respectively. On SAKURA, we observe a nearly 10% increase in overall accuracy, along with an 8% improvement specifically on multi-hop questions. In contrast, improvements in Audio Entailment are relatively smaller. Overall, we find that BALSa-MA, which enables the model to learn from two audio inputs simultaneously, enhances its basic audio understanding abilities, thereby benefiting more advanced reasoning tasks.
@@
VII Conclusion
In this work, we introduce BALSa, a framework that synthesizes general-purpose caption-style alignment data for training ALLM. This framework reduces the reliance on large collections of task-specific QA datasets. Our experimental results demonstrate that the proposed approach achieves comparable or superior performance to previous approaches while improving data and computational efficiency. BALSa effectively mitigates audio hallucinations, further reinforcing its robustness in real-world applications. Beyond single-audio alignment, we extend to more complex audio understanding settings. In multi-audio scenarios, this extension further enhances modality alignment and leads to better performance on audio understanding and reasoning tasks. We further examine the role of discrimination-based training and observe that learning to handle multiple audio inputs alone is sufficient to improve overall understanding, without requiring explicit comparative objectives. Overall, this paper provides a scalable and efficient pathway for developing and advancing audio-aware large language models.